# Examining Post-Training Quantization for Mixture-of-Experts: A Benchmark

### v1

@@ 4.2 Benchmark Results
Expert usage frequency is a popular heuristic in the compression of MoE models, predicated on the insight that less frequently used experts are likely less crucial. Our experiments, detailed in Table 1, corroborate its effectiveness as a quantization heuristic for MoE models. In particular, for the DeepSeek-MoE-16B-base model, this heuristic markedly outperforms the strategy of randomly allocating more bits to experts, likely due to the model’s unbalanced routing distribution. However, with the Mixtral-8x7B model, where the routing distribution is more balanced, the advantage of using expert usage frequency over random allocation is less significant.
Table 1: Comparison of the expert usage frequency heuristic v.s. random allocation. For the Mixtral-8x7B model, we compare the allocation of $4$ bits to the top-{$2$, $4$} most frequently used experts per MoE block against randomly selecting {$2$, $4$} experts for the same bit allocation. For the DeepSeek-MoE-16B-base model, we keep shared expert {$8$} bits and compare between top-{$10$, $15$, $20$, $25$} most frequently used experts against randomly selecting {$10$, $15$, $20$, $25$} experts per MoE block. The remaining experts are quantized to $2$ bits, while all attention layers are uniformly quantized to $4$ bits. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ independent trials.
@@
Figure 2: Comparison of quantizing more bits for attention vs. FFNN. It is evaluated on the Mixtral-8x7B model. FFNN results show the mean and standard deviation (error bars) from $3$ independent trials.
@@
Because of the unique characteristics of the feedforward neural network (FFNN) within the mixture of experts (MoE) framework. we explore the attention layer and the feedforward neural network layer, which deserves more bits. We compare the performance evaluated by quantizing the attention layers with more bits v.s. randomly selecting experts in the FFNN layers with more bits, maintaining the same average bits of the entire MoE model for a fair comparison. Specifically, we quantize the attention weight or randomly selected FFNN weight to {$2$, $4$, $8$} bits, while All other weights are quantized to $2$ bits by default. As illustrated in Figure 2, quantizing attention weights to higher bit levels (i.e., $4$ or $8$ bits) consistently results in significant performance gains (over $5\%$) under each average bit allocation for the MoE model. This greater efficiency likely stems from the fact that attention weights are activated for every token, while FFNN weights only engage with a subset of the input tokens. Consequently, increasing the quantization bits for FFNN weights does not benefit all inputs. Based on these findings, attention weights are quantized to $4$ bits by default in all following experiments.
Table 2: Comparison between quantizing first $k$ v.s. last $k$ MoE blocks with higher (i.e. $4$) bits. All weights in attention layers are quantized to $4$ bits, and the other weights are quantized to $2$ bits. In DeepSeek-MoE-16B-base model, we keep the first block that is dense block as $4$ bits by default. We evaluate $k$ of $4$ and $8$. The higher performance of each comparison pair is marked as bold.
@@
As more and more Mixture-of-Experts (MoE) architectures emerge, we investigate which layer of the MoE block is more critical and thus deserves more bits during the quantization process. As shown in Table 2, we evaluate the performance of allocating more bits to the first $k$ blocks versus the last $k$ blocks in quantization. The results consistently indicate that higher bit quantization of the first few blocks yields better performance, suggesting that we can allocate more bits to the quantization of the first blocks of the model. This observation aligns with prior studies that have empirically confirmed the greater importance of the first few Transformer blocks [6, 20].
Figure 3: Comparison of quantizing more bits for shared experts vs. others experts. "Others" results show the mean and standard deviation from $3$ independent trials of random selecting $2$ experts from the non-shared experts.
@@
The DeepSeek-MoE-16B-base model includes two shared experts within each MoE block to obtain common knowledge across varying domains and alleviate the parameter redundancy. To evaluate their role in quantization, we compare quantizing these two shared experts with more bits v.s. randomly selecting two non-shared experts for more bit allocation, maintaining the same average bits for a fair comparison. The shared or random non-shared experts are quantized to $2$, $4$, $8$ bits, while attention weights are set to $4$ bits and all other weights to $2$ bits. As depicted in Figure 3, allocating higher bit levels (i.e., $4$ or $8$ bits) to shared experts consistently yields superior performance. This enhanced efficiency and effectiveness are attributed to the shared experts being activated for every input token, unlike non-shared experts, which only engage with specific subsets of the tokens. Allocating more quantization bits to shared experts thus proves to be both more efficient and effective.
5 Extended Study to Improve MoE Quantization
@@
In this section, we expand our benchmark results from weight quantization to include activation quantization. Additionally, we introduce two novel algorithmic advancements aimed at enhancing the effectiveness of identifying crucial components within MoE models for improved quantization performance.
5.1 Quantizing Both Weight and Activation
We further expand our study by simultaneously including weight and activation quantization to validate our conclusions. Specifically, we employ SmoothQuant [30] combined with our expert-usage-frequency heuristic. It selects the top-$2$ experts’ weights per MoE block in the Mixtral-8x7B model and the top-$16$ experts’ weights per MoE block in the DeepSeek-MoE-16B-base for quantization to $4$ bits, while quantizing all other weights to $2$ bits. The evaluation results, presented in Table 3, reveal the marginal performance gap across different activation quantization bits. This demonstrates that our conclusions regarding weight quantization are robust and can be reliably extended to various activation quantization scenarios as well.
Table 3: Combination of activation quantization with the expert-usage-based heuristic. We evaluate it on the top-$2$ most frequently used experts per MoE block in Mixtral-8x7B and the top-$16$ frequent experts per MoE block in DeepSeek-MoE-16B-base, quantizing these experts to $4$ bits. All attention weights are also quantized to $4$ bits, while all other weights are quantized to $2$ bits. The higher performance of each comparison pair is marked as bold.
Weight Bits Activation Bits WinoGrande (%) COPA (%) OBQA (%) HellaSwag (%) PIQA (%) MMLU (%) Average (%)
Mixtral-8x7B
$2.54$ $4$ $\mathbf{50.28}$ $51.00$ $\mathbf{26.80}$ $25.99$ $\mathbf{51.90}$ $23.85$ $38.30$
$8$ $50.04$ $\mathbf{60.00}$ $\mathbf{26.80}$ $\mathbf{26.55}$ $51.58$ $23.77$ $\mathbf{39.79}$
$16$ $49.41$ $\mathbf{60.00}$ $26.60$ $26.53$ $51.85$ $\mathbf{23.86}$ $39.71$
DeepSeek-MoE-16B-base
$2.71$ $4$ $48.22$ $\mathbf{53.00}$ $27.20$ $26.12$ $50.65$ $\mathbf{26.86}$ $38.67$
$8$ $49.96$ $51.00$ $\mathbf{27.60}$ $\mathbf{26.58}$ $\mathbf{53.86}$ $25.91$ $\mathbf{39.15}$
$16$ $\mathbf{50.19}$ $51.00$ $\mathbf{27.60}$ $26.43$ $53.70$ $25.16$ $39.01$
5.2 Concentrating Linear Layers with Larger Weight Outliers
Insight.
@@ 5.2 Concentrating Linear Layers with Larger Weight Outliers
From the quantization perspective, the larger the range of a weight magnitude group, the more difficult it will be for quantization. We found that, in MoE, each FFNN linear weight matrix consists predominantly of values within a narrow range, interspersed with a few significant outliers. Consequently, we propose a weight-magnitude-based metric to identify those linear layers that are challenging to quantize effectively, thereby necessitating a higher allocation of quantization bits.
Methodology.
@@
Experiments.
We evaluate this metric by comparing its application for the top-$p\%$ of linear layers against randomly selecting linear layers, using percentages of $25\%$ and $50\%$. In DeepSeek-MoE-16B-base model, we also involve shared experts using this metric. As illustrated in Table 4, our proposed scorer consistently outperforms the random baseline on both models and almost all tasks (except HellaSwag and MMLU). This is particularly evident in the DeepSeek-MoE-16B-base model, where it achieves an average performance improvement of about $3\%$, aligning with our expectations.
Table 4: Comparison between using our linear weight scorer vs. random selection of linear layers for bit allocation in quantization. We evaluate by quantizing $25\%$ of the linear layers across all MoE blocks (i.e., FFNN) to $4$ bits. All attention weights are quantized to $4$ bits, and all other weights are quantized to $2$ bits. In each comparison pair, the higher performance is highlighted in bold. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ independent trials.
@@
Visualization.
As shown in Figure 4, we visualize the proposed outlier-score for each FFNN linear weight within the Mixtral-8x7B model. Given that each FFNN expert includes three linear layers, namely the gate projection, up projection, and down projection, we visualize these components separately to ensure clarity. Notably, many of the down projection linear layers, particularly those positioned later in the MoE model, exhibit significantly higher outlier-scores compared to others.
Figure 4: Visualization of the outlier-score metric applied to each FFNN linear weight matrix within the Mixtral-8x7B model. For clearer visualization, we present separate components, including the gate projection (left), up projection (middle), and down projection (right) in FFNN experts.
5.3 Training Block Quantization Importance Score Predictor.
Insight.
We find an increasing cosine similarity between the tensors generated before and after the FFN blocks for some of the MoE blocks, indicating less important computation results produced by these blocks. This observation also aligns with observations on dense models in previous literature [12]. Therefore, the basic idea is that less accurate output of these blocks producing tokens with high cosine similarity will not affect the overall model performance much, thus lower weight bits might not hurt performance much.
Methodology.
To capture the generalized hidden states’ dynamic information of each MoE block, we train a small two-layer FFNN with a tangent activation function. This network predicts the cosine similarity between the input and output hidden states. We utilize a dataset of $400$ random sequences, each containing $1024$ tokens from the WikiText dataset [21], for training. The detailed training procedure is in Algorithm 2. During quantization, we employ this predictor to run inference on the calibration data, computing the average predicted score for each MoE block across all tokens. A higher predicted score indicates less important and fewer bits for quantization.
Algorithm 2 The Training Procedure of Block Score Predictor.
2:Let $\mathcal{BSP}$ denotes the block score predictor.
3:$\mathcal{X}\leftarrow\left\{\mathbf{x}_{i}\mid i\in[N]\right\}$
5:$\mathcal{BSP}\leftarrow\texttt{train}(\mathcal{X},\mathcal{S})$
6:Return: The importance score predictor $\mathcal{BSP}$ for MoE Block $M$.
Experiments.
In Table 5, we compare the performance of using our block importance predictor to select $k$ MoE blocks for $4$ bits and others for $2$ bits quantization with two other baselines: ① random selecting $k$ MoE blocks, and ② first $k$ MoE blocks (as it is the best in Q3 in Section 4.2). Evaluation results on the DeepSeek-MoE-16B-base model are presented in Table 5, showing the superiority of our method against the other two baselines.
Table 5: Comparison between using our MoE block importance predictor v.s. two baselines: ①random selecting and ②first $k$ MoE blocks. The predicted or selected MoE blocks are quantized to $4$ bits, all attention weights are quantized to $4$ bits, and all other weights are quantized to $2$ bits. In each comparison, the highest performance is highlighted in bold. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ independent trials.
@@
Figure 5: Visualization of the predicted MoE block importance score using our trained predictors.
Visualization.
We visualize the predicted scores of each MoE block using our trained predictors in the DeepSeek-MoE-16B-base model, as shown in Figure 5. Notably, MoE blocks situated in the middle of the model, which exhibit higher scores, are regarded as less critical. Consequently, these blocks will be quantized with fewer bits (specifically, $2$ bits), reflecting their lower importance. Besides, Figure 5 also demonstrates that the first few MoE blocks are more important aligned with Q3. Interestingly, the last two blocks of the DeepSeek-MoE-16B-base model are also crucial, thereby allocating more bits and yielding better performance.
6 Conclusion
This work investigates various heuristic-based MoE quantization methods in the post-training setting. While vanilla quantization techniques (e.g., GPTQ) prove less effective and efficient when applied directly to MoE models, determining which MoE model components should be allocated more quantization bits remains an open question. We present the first benchmark study on MoE quantization, revealing critical heuristic-based principles, such as the importance disparities among different MoE blocks. Drawing on these insights, we introduce innovative techniques, including a block importance predictor and a linear layer outlier range scorer, to more precisely identify components that benefit from increased bit quantization. These methods substantially improve the quantization process’s effectiveness and efficiency for MoE models.
@@ A.1 Evaluation Datasets
WinoGrande [1] is a large-scale dataset designed for commonsense reasoning, consisting of pronoun resolution problems. Each instance in the dataset presents a sentence with an ambiguous pronoun that needs to be resolved based on context. This task tests the model’s ability to understand and reason about everyday situations.
The Choice of Plausible Alternatives (COPA) dataset [10] focuses on causal reasoning. Each question in COPA consists of a premise and two choices, where the model must select the more plausible alternative. This task evaluates the model’s understanding of cause-and-effect relationships in natural language.
OpenBookQA [22] is a multiple-choice question-answering dataset that requires the model to use both scientific facts and commonsense knowledge. The dataset challenges the model’s ability to combine factual knowledge with reasoning to answer questions correctly.
HellaSwag [31] is a benchmark for commonsense NLI (Natural Language Inference) that tests the model’s ability to predict the most plausible continuation of a given sentence. The dataset contains scenarios from various domains, such as cooking and sports, requiring the model to understand context and plausibility.
The Massive Multitask Language Understanding (MMLU) benchmark [11] evaluates models across a wide range of subjects, from elementary mathematics to law. For this study, we report performance on MMLU with a 5-shot setting, where the model is given five examples per task before evaluation, allowing us to gauge the model’s few-shot learning capabilities.

### v2

@@ 4.3. Benchmark Results
Expert usage frequency is a popular heuristic in the compression of MoE models, predicated on the insight that less frequently used experts are likely less crucial. Our experiments, detailed in Table 2, corroborate its effectiveness as a quantization heuristic for MoE models. In particular, for the DeepSeek-MoE-16B-base model, this heuristic outperforms the strategy of randomly allocating more bits to experts, likely due to the model’s unbalanced routing distribution. However, with the Mixtral-8x7B model, where the routing distribution is more balanced, the advantage of using expert usage frequency over random allocation is less significant.
Table 2. Comparison of the expert usage frequency heuristic v.s. random allocation. For the Mixtral-8x7B model, we compare the allocation of $4$ bits to the top-{$2$, $4$} most frequently used experts per MoE block against randomly selecting {$2$, $4$} experts for the same bit allocation. For the DeepSeek-MoE-16B-base model, we keep shared expert {$8$} bits and compare between top-{$10$, $15$, $20$, $25$} most frequently used experts against randomly selecting {$10$, $15$, $20$, $25$} experts per MoE block. The remaining experts are quantized to $2$ bits, while all attention layers are uniformly quantized to $4$ bits. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ independent trials.
@@
Figure 4. Comparison of quantizing more bits for attention vs. FFNN and shared experts v.s. others, evaluated on the Mixtral-8x7B model. FFNN and others’ results show the mean and standard deviation (error bars) from $3$ independent trials.
@@
Because of the unique characteristics of the feedforward neural network (FFNN) within the mixture of experts (MoE) framework. we explore the attention layer and the feedforward neural network layer, which deserves more bits. We compare the performance evaluated by quantizing the attention layers with more bits v.s. randomly selecting experts in the FFNN layers with more bits, maintaining the same average bits of the entire MoE model for a fair comparison. Specifically, we quantize the attention weight or randomly selected FFNN weight to {$2$, $4$, $8$} bits, while All other weights are quantized to $2$ bits by default. As illustrated in Figure 4 (left), quantizing attention weights to higher bit levels (i.e., $4$ or $8$ bits) consistently results in significant performance gains (over $5\%$) under each average bit allocation for the MoE model. This greater efficiency likely stems from the fact that attention weights are activated for every token, while FFNN weights only engage with a subset of the input tokens. Consequently, increasing the quantization bits for FFNN weights does not benefit all inputs. Based on these findings, attention weights are quantized to $4$ bits by default in all following experiments.
Table 3. Comparison between quantizing first $k$ v.s. last $k$ MoE blocks with higher (i.e. $4$) bits. All weights in attention layers are quantized to $4$ bits, and the other weights are quantized to $2$ bits. In DeepSeek-MoE-16B-base model, we keep the first block that is dense block as $4$ bits by default. We evaluate $k$ of $4$ and $8$. The higher performance of each comparison pair is marked as bold.
@@
As more and more Mixture-of-Experts (MoE) architectures emerge, we investigate which layer of the MoE block is more critical and thus deserves more bits during the quantization process. As shown in Table 3, we evaluate the performance of allocating more bits to the first $k$ blocks versus the last $k$ blocks in quantization. The results consistently indicate that higher bit quantization of the first few blocks yields better performance, suggesting that we can allocate more bits to the quantization of the first blocks of the model. This observation aligns with prior studies that have empirically confirmed the greater importance of the first few Transformer blocks (Dai et al., 2024; Ma et al., 2023).
@@
The DeepSeek-MoE-16B-base model includes two shared experts within each MoE block to obtain common knowledge across varying domains and alleviate the parameter redundancy. To evaluate their role in quantization, we compare quantizing these two shared experts with more bits v.s. randomly selecting two non-shared experts for more bit allocation, maintaining the same average bits for a fair comparison. The shared or random non-shared experts are quantized to $2$, $4$, $8$ bits, while attention weights are set to $4$ bits and all other weights to $2$ bits. As depicted in Figure 4 (right), allocating higher bit levels (i.e., $4$ or $8$ bits) to shared experts consistently yields superior performance. This enhanced efficiency and effectiveness are attributed to the shared experts being activated for every input token, unlike non-shared experts, which only engage with specific subsets of the tokens. Allocating more quantization bits to shared experts thus proves to be both more efficient and effective.
5. Extended Study to Improve MoE Quantization
@@
In this section, we introduce two novel data-driven techniques aimed at enhancing the effectiveness of identifying crucial components within MoE models for improved quantization performance.
Figure 5. (a) Visualization of the outlier-aware linear layer scorer metric applied to each FFNN linear weight matrix within the Mixtral-8x7B model. For clearer visualization, we present separate components, including the gate projection (left), up projection (middle), and down projection (right) in FFNN experts. (b) Visualization of the MoE block importance score predictor metric applied on the DeepSeek-MoE-16B-base model.
5.1. Outlier-Aware Linear Layer Scorer
@@
Insight. From the quantization perspective, the larger the range of a weight magnitude group, the more difficult it will be for quantization. We found that, in MoE, each FFNN linear weight matrix consists predominantly of values within a narrow range, interspersed with a few significant outliers. Consequently, we propose a weight-magnitude-based metric to identify those linear layers that are challenging to quantize effectively, thereby necessitating a higher allocation of quantization bits.
@@
Experiments. We evaluate this metric by comparing its application for the top-$p\%$ of linear layers against randomly selecting linear layers, using percentages of $25\%$ and $50\%$. In DeepSeek-MoE-16B-base model, we also involve shared experts using this metric. As illustrated in Table 4, our proposed scorer consistently outperforms the random baseline on both models and almost all tasks (except HellaSwag and MMLU). This is particularly evident in the DeepSeek-MoE-16B-base model, where it achieves an average performance improvement of about $3\%$, aligning with our expectations.
Table 4. Comparison between using our linear weight scorer vs. random selection of linear layers for bit allocation in quantization. We evaluate by quantizing $25\%$ of the linear layers across all MoE blocks (i.e., FFNN) to $4$ bits. All attention weights are quantized to $4$ bits, and all other weights are quantized to $2$ bits. In each comparison pair, the higher performance is highlighted in bold. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ trials.
@@
Table 5. Comparison between using our MoE block importance predictor v.s. two baselines: ①random selecting and ②first $k$ MoE blocks. The predicted or selected MoE blocks are quantized to $4$ bits, all attention weights are quantized to $4$ bits, and all other weights are quantized to $2$ bits. In each comparison, the highest performance is highlighted in bold. All random experimental results in the format of $a\pm b$ provide the mean value $a$ and its standard deviation $b$ over $3$ independent trials.
@@
Visualization. As shown in Figure 5 (a), we visualize the proposed outlier-score for each FFNN linear weight within the Mixtral-8x7B model. Given that each FFNN expert includes three linear layers, namely the gate projection, up projection, and down projection, we visualize these components separately to ensure clarity. Notably, many of the down projection linear layers, particularly those positioned later in the MoE model, exhibit significantly higher outlier-scores compared to others.
5.2. MoE Block Importance Score Predictor
Algorithm 2 The Training Procedure of Block Score Predictor.
2:Let $\mathcal{BSP}$ denotes the block score predictor.
3:$\mathcal{X}\leftarrow\left\{\mathbf{x}_{i}\mid i\in[N]\right\}$
5:$\mathcal{BSP}\leftarrow\texttt{train}(\mathcal{X},\mathcal{S})$
6:Return: The importance scorer $\mathcal{BSP}$ for MoE Block $M$.
Insight. We find an increasing cosine similarity between the tensors generated before and after the FFN blocks for some of the MoE blocks, indicating less important computation results produced by these blocks. This observation also aligns with observations on dense models in previous literature (Jaiswal et al., 2024). Therefore, the basic idea is that less accurate output of these blocks producing tokens with high cosine similarity will not affect the overall model performance much, thus lower weight bits might not hurt performance much.
Methodology. To capture the generalized hidden states’ dynamic information of each MoE block, we train a two-layer FFNN with a tangent activation function. This network predicts the cosine similarity between the input and output hidden states. We utilize a dataset of $400$ random sequences, each containing $1024$ tokens from the WikiText dataset (Merity et al., 2016), for training. The detailed training procedure is in Algorithm 2. During quantization, we applied this predictor to the calibration data, computing an average predicted score for each MoE layer across all tokens. A higher predicted score indicates less importance and fewer bits for quantization.
Experiments. In Table 5, we compare the performance of using our block importance predictor to select $k$ MoE blocks for $4$ bits and others for $2$ bits quantization with two other baselines: ① random selecting $k$ MoE blocks, and ② first $k$ MoE blocks (as it is the best in Q3 in Section 4.3). Evaluation results on the DeepSeek-MoE-16B-base model are presented in Table 5, showing the superiority of our method against the other two baselines.
Visualization. We visualize the predicted scores of each MoE block using our trained predictors in the DeepSeek-MoE-16B-base model, as shown in Figure 5 (b). Notably, MoE blocks situated in the middle of the model, which exhibit higher scores, are regarded as less critical. Consequently, these blocks will be quantized with fewer bits (specifically, $2$ bits), reflecting their lower importance. Besides, Figure 5 (b) also demonstrates that the first few MoE blocks are more important aligned with Q3. Interestingly, the last two blocks of the DeepSeek-MoE-16B-base model are also crucial, thereby allocating more bits and yielding better performance.
6. Conclusion
This work introduces mixed-precision MoE quantization through a systematic investigation of various heuristic-based approaches in the post-training setting. While conventional quantization techniques (e.g., GPTQ) show limited effectiveness when directly applied to MoE models, the question of optimal bit allocation across different MoE model components requires deeper exploration. We present QuantMoE-Bench, the first comprehensive benchmark to comprehensively study post-training quantization of MoE models, which reveals critical insights, including significant importance variations among MoE blocks. Drawing on these insights, we further develop innovative techniques, i.e., a block importance predictor and a linear layer outlier range scorer, to more precisely identify which components benefit most from higher-precision quantization. Our methods substantially improve the effectiveness of MoE model quantization. We hope our study provides novel insights into the memory compression of MoE-based models.
@@ A.1. Evaluation Datasets
WinoGrande (ai2, 2019) is a large-scale dataset designed for commonsense reasoning, consisting of pronoun resolution problems. Each instance in the dataset presents a sentence with an ambiguous pronoun that needs to be resolved based on context. This task tests the model’s ability to understand and reason about everyday situations.
The Choice of Plausible Alternatives (COPA) dataset (Gordon et al., 2012) focuses on causal reasoning. Each question in COPA consists of a premise and two choices, where the model must select the more plausible alternative. This task evaluates the model’s understanding of cause-and-effect relationships in natural language.
OpenBookQA (Mihaylov et al., 2018) is a multiple-choice question-answering dataset that requires the model to use both scientific facts and commonsense knowledge. The dataset challenges the model’s ability to combine factual knowledge with reasoning to answer questions correctly.
HellaSwag (Zellers et al., 2019) is a benchmark for commonsense NLI (Natural Language Inference) that tests the model’s ability to predict the most plausible continuation of a given sentence. The dataset contains scenarios from various domains, such as cooking and sports, requiring the model to understand context and plausibility.
The Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021) evaluates models across a wide range of subjects, from elementary mathematics to law. For this study, we report performance on MMLU with a 5-shot setting, where the model is given five examples per task before evaluation, allowing us to gauge the model’s few-shot learning capabilities.